# zkFuzz: Foundation and Framework for Effective Fuzzing of Zero-Knowledge Circuits

### v1

@@ 2.3. Circom Primer
[…]
@@
Equality Constraint ($=$) asserts equality between two signals and generates a constraint. If the constraint_ assert_ disabled flag is set, the assertion is not inserted into the computation.
@@
Other than $=$, Circom operators do not allow adding constraints without performing the corresponding computation, so $=$ with constraint_assert_disabled set is the only way to introduce over-constrained bugs. Nonetheless, over-constrained circuits can arise in other languages such as halo2 (Zcash, 2022; Soureshjani et al., 2023), making it important to define them formally.
@@
3. Definitions of ZK Program Vulnerabilities
This section formulates the typical ZK circuit’s vulnerabilities as the Trace-Constraint Consistency Test (TCCT), a unified model that rigorously formulates bugs within ZK circuits as the inconsistency between the computation logic and its associated circuit constraint.
3.1. Incompleteness of Prior Definitions
Prior formal definitions of ZK program vulnerabilities are incomplete on two fronts. First, when detecting under-constrained vulnerabilities, they focus on identifying nondeterministic circuit constraints (Pailoor et al., 2023; Jiang et al., 2025), while ignoring computation aborts. If the circuit constraints $\mathcal{C}$ allow the same input $x$ to map to an output $y^{\prime}$ different from the computed output $y$, the circuit is clearly under-constrained. However, even if $\mathcal{C}$ is deterministic—accepting only one value of $y$ for a given $x$—the circuit can still be under-constrained if the computation $\mathcal{P}$ aborts on $x$ (e.g., due to a runtime assertion). Attackers can exploit such aborts by crafting inputs that cause the computation to fail while still satisfying the circuit constraints, thereby producing bogus proofs for traces that never actually occurred. Our evaluation shows that such abort vulnerabilities account for nearly 50% of the real-world bugs zkFuzz identified.
Second, when detecting over-constrained vulnerabilities, prior work (Chen et al., 2024) considers only a special case where $\mathcal{C}$ permits an empty set of traces. A complete definition must consider the full set of execution traces produced by $\mathcal{P}$ and those accepted by $\mathcal{C}$. In particular, it must account for intermediate values. Suppose $\mathcal{P}$ produces a trace $(x,z,y)$ for input $x$, but $\mathcal{C}$ accepts only $(x,z^{\prime},y)$ where $z\neq z^{\prime}$. In this case, $\mathcal{C}$ is over-constrained with respect to $\mathcal{P}$, as an honest prover with the valid trace $(x,z,y)$ would be unable to generate a valid proof. If, however, $\mathcal{C}$ accepts both $(x,z,y)$ and $(x,z^{\prime},y)$ (a common optimization in practice), there is no violation, as the prover can produce a valid proof.
3.2. Trace-Constraint Consistency Test
Let $\mathbb{F}_{q}$ be a finite field of order $q$, where $q$ is a prime number.
Definition 3.1 (ZK Program).
A ZK program is defined as a pair of a computation and a set of circuit constraints $(\mathcal{P},\mathcal{C})$, where:
@@ 3.2. Trace-Constraint Consistency Test
  • On abnormal termination, it returns the special value (_,)_bottom(\_,$\bot$)( _ , ⊥ ).
@@
Definition 3.2 (Execution Trace).
@@
Definition 3.4 (Constraint Satisfaction Set).
@@
$\mathcal{S}(\mathcal{C}):=\{(x,z,y)\mid x\in\mathbb{F}^{n}_{q},z\in\mathbb{F}^{k}_{q},y\in\mathbb{F}^{m}_{q},C(x,z,y)=\text{true}\}$
@@
Definition 3.5 (Projection).
[…]
@@
Definition 3.6 (Under-Constrained Circuit).
@@
(1) $\Pi_{xy}(\mathcal{S}(\mathcal{C}))\setminus\Pi_{xy}(\mathcal{T}(\mathcal{P}))\neq\emptyset$
Intuitively, $\mathcal{C}$ is under-constrained for $\mathcal{P}$ if $\mathcal{C}$ accepts $(x,\_,y)$ but $\mathcal{P}(x)=(\_,y^{\prime})$ and $y\neq y^{\prime}$ or $\mathcal{P}(x)$ aborts. Projection out the intermediate values here is necessary for the following reason. In contrast to over-constrained vulnerabilities, if $\mathcal{P}(x)=(z,y)$ and $\mathcal{C}$ accepts $(x,z^{\prime},y)$ such that $z\neq z^{\prime}$, it is not an under-constrained vulnerability, since malicious provers still cannot create bogus proofs asserting that $\mathcal{P}(x)$ yields an output $y^{\prime}\neq y$. Similarly, even if $\mathcal{C}$ accepts both $(x,z,y)$ and $(x,z^{\prime},y)$, adversaries cannot forge proofs with outputs differing from $y$.
This definition naturally takes care of abnormal termination of the computation. Any input $x$ that causes $\mathcal{P}$ to abort will not appear in the trace set $\mathcal{T}(\mathcal{P})$. If $\mathcal{S}(\mathcal{C})$ includes any tuple of the form $(x,\_,\_)$, then the circuit is under-constrained, as it accepts an input that the computation cannot successfully process.
Definition 3.7 (Over-Constrained Circuit).
@@
(2) $\mathcal{T}(\mathcal{P})\setminus\mathcal{S}(\mathcal{C})\neq\emptyset$
@@
We now combine Eq. 1 and Eq. 2 and define a test for the consistency between the trace and the constraint satisfaction sets:
Definition 3.8 (Trace-Constraint Consistency Test).
@@
(3) $\Pi_{xy}(\mathcal{S}(\mathcal{C}))\setminus\Pi_{xy}(\mathcal{T}(\mathcal{P}))=\emptyset\hskip 5.0pt\land\hskip 5.0pt\mathcal{T}(\mathcal{P})\setminus\mathcal{S}(\mathcal{C})=\emptyset$
If Eq. 3 holds, we say that $\langle\mathcal{P},\mathcal{C}\rangle$ is well-constrained. This condition is equivalent to $\mathcal{T}(\mathcal{P})\subseteq\mathcal{S}(\mathcal{C})$land $\Pi_{xy}(\mathcal{T}(\mathcal{P}))=\Pi_{xy}(\mathcal{S}(\mathcal{C}))$, meaning that all execution traces satisfy the constraints, and the set of input-output pairs is identical between the execution traces and the constraint satisfaction set.
@@
a) Determining whether $\mathcal{C}$ is under-constrained for $\mathcal{P}$ is NP-complete.
b) Determining whether $\mathcal{C}$ is over-constrained for $\mathcal{P}$ is NP-complete.
c) The Trace-Constraint Consistency Test is co-NP-complete.
Proof is based on the reduction from the Boolean Satisfiability problem (SAT) and the Boolean tautology problem and can be found in Appendix A.
Comparison to Prior Definitions. Tab. 1 compares our model, TCCT, with prior definitions. For under-constrained bugs, prior work detect only non-deterministic behaviors (Pailoor et al., 2023; Jiang et al., 2025), and fail to capture those caused by computation aborts. For over-constrained bugs, prior work considers only a special case where the constraint satisfaction set is empty (Chen et al., 2024). In contrast, TCCT is more general and complete, capturing all bugs detectable by existing models, as well as additional bugs that prior definitions overlook. A key advantage of TCCT is its DSL-independence: even though different DSLs may offer varied primitives or syntactical sugar for specifying constraints and computations in ZK circuits, TCCT ’s definitions consistently apply across all of them.
Table 1. Comparison of formal definitions of ZK vulnerabilities. Our TCCT is more general than existing models.
@@
(Chen et al., 2024) Same as (Pailoor et al., 2023; Jiang et al., 2025) $\mathcal{S}(\mathcal{C})=\emptyset$
@@
3.3. Vulnerable and Benign Examples
@@
1template Verify() {
2 signal input x1;
3 signal input x2;
4
5 component h = Hash();
6 h.x <== x1;
7 assert(x2==h.y);
@@
Code 4: Under-constrained circuit from (Aurobindo, 2025) undetectable by existing formal definitions. Only our TCCT can capture it, and zkFuzz successfully identifies it.
@@
8 x*y $=$ 0;
9}
Code 5: Benign circuit from (iden3, 2025) incorrectly flagged by prior static tools. Its trace and constraint satisfaction sets are shown in Tab. 2.
Code 4 shows a real-world under-constrained circuit (slightly modified for clarity) caught by zkFuzz from circom-zkVerify (Aurobindo, 2025). it had previously gone undetected until our discovery, later confirmed and fixed by the developers. The code instantiates a Hash template as component h (line 5), computes the hash value of input x1 (line 6), and checks whether it matches x2 using assert (line 7). Circom’s assert, however, performs runtime checks only without adding any constraint. As a result, the computation aborts when x1 does not hash to x2, the constraints remain satisfiable for any combination of x1 and x2. No prior formal tools can detect this bug because the circuit has no output and is thus always deterministic. While a static tool (Wen et al., 2024) does flag this code, it relies on pattern matching and produces many false positives. The fact that this bug went unfixed indicates that suggests that developers who may have run the tool did not inspect all of its warnings. To fix this circuit, assert should be replaced with $=$, which explicitly encodes the equality check into the constraints.
Table 2. Trace and Constraint Satisfaction Sets of IsZero in Code 5 ($q=3$)
@@
Code 5 shows a benign circuit template from Circom’s de facto standard library, circomlib (iden3, 2025), which returns one if the input $x$ is zero and zero otherwise. Line 6 introduces an intermediate value $z$ using a weak assignment, which does not add any constraint since it is not a quadratic expression. Line 7 assigns the output $y$ and introduces a constraint relating $x$, $z$, and $y$. Line 8 adds both a runtime assertion and an equality constraint. The full execution trace and constraint satisfaction sets are shown in Tab. 2. Notably, $\mathcal{T}(\mathcal{P})$ and $\mathcal{S}(\mathcal{C})$ are not identical: $\mathcal{S}(\mathcal{C})$ includes two additional tuples. However, TCCT correctly handles this case. After projecting out intermediate values, the sets of $(x,y)$ pairs are identical, indicating no under-constrained bug. While $\mathcal{S}(\mathcal{C})$ includes three different intermediate values, $z=0,1,2$, for the case where $x=0$ and $y=1$, it includes the valid tuple $(0,0,1)$ from $\mathcal{T}(\mathcal{P})$, so there is no over-constrained bug either. Despite being benign, this circuit is incorrectly flagged by prior static analyzers, which rely on pattern matching rather than a precise semantic definition of ZK bugs.
More examples, including under-constrained circuits caused by insecure use of the LessThan template, a bug that only zkFuzz can detect; and over-constrained circuits can be found in Appendix B.
4. zkFuzz Design
@@
4.1. Joint Program Mutation & Input Generation
zkFuzz uses both program mutation and input fuzzing for detecting vulnerabilities. Although program mutation is widely used in mutation testing to assess test case effectiveness (Papadakis et al., 2019), our objective differs: we leverage mutation specifically to generate alternative traces that continue to satisfy the constraint, exposing undesired behaviors in ZK circuits.
Algo. 1 describes the key steps that examine both $\mathcal{P}$ and its associated circuit constraints $\mathcal{C}$ by using two testing modes:
@@ 4.1. Joint Program Mutation & Input Generation
Algorithm 1 Fuzzing with Program Mutation to solve TCCT
@@
For simplicity, Algo. 1 shows one input and one program mutant sampled per iteration. In practice, we set the number of input samples and program mutants to 30. Both input sampling and program mutation can be biased toward targets that are more likely to expose vulnerabilities. The algorithm also requires minimal bookkeeping across iterations: once program mutants are generated, the current iteration operates independently of the previous ones.
Figure 2. Basic workflow of zkFuzz. This fuzzing framework systematically mutates program logic and feeds artificially generated input data to the original and mutated programs to catch inconsistencies between the program and the constraint. The error-based fitness function is utilized to steer the selection and crossover of mutants. Several target selectors are also applied for input generation and program mutation to guide the search.
4.2. Guided Search

### v2

@@ 2.3 Circom Primer
Value Domains and Operators. In Circom, all computations are performed over a finite field $\mathbb{F}_{q}$, where $q$ is a large prime number. The language supports standard arithmetic operators such as addition (++), subtraction (-), multiplication (*), and division (//), all performed modulo $q$. Additionally, Circom provides integer division ($\backslash$), bitwise, and comparison operators for defining complex circuit logic. The operands or results of these operators often need implicit conversion between integer and finite fields.
@@
Equality Constraint (===) asserts equality between two signals and generates a constraint. If the constraint_ assert_ disabled flag is set, the assertion is not inserted into the computation.
@@
Other than ===, Circom operators do not allow adding constraints without performing the corresponding computation, so === with constraint_assert_disabled set is the only way to introduce over-constrained bugs. Nonetheless, over-constrained circuits can arise in other languages such as zigren [19] and halo2 [41, 42], making it important to define them formally.
@@
3 Definitions of ZK Program Vulnerabilities
This section formulates the typical ZK circuit’s vulnerabilities as the Trace-Constraint Consistency Test (TCCT), a unified language-agnostic model that rigorously formulates bugs within ZK circuits as discrepancies between traces produced by a witness-generation program on possible inputs and the constraints enforced by the circuit.
3.1 Incompleteness of Prior Definitions
Prior formal definitions of ZK program vulnerabilities are incomplete on two key fronts, as we elaborate below. Both gaps stem from prior work’s exclusive focus on constraints while neglecting the computation logic in witness generation.
First, when detecting under-constrained vulnerabilities, they focus on identifying nondeterministic circuit constraints [18, 25], while ignoring computation aborts. If the circuit constraints $\mathcal{C}$ allow the same input $x$ to map to an output $y^{\prime}$ different from the computed output $y$, the circuit is clearly under-constrained. However, even if $\mathcal{C}$ is deterministic—accepting only one value of $y$ for a given xx—the circuit can still be under-constrained if the computation $\mathcal{P}$ aborts on $x$ (e.g., due to a runtime assertion). Attackers can exploit such aborts by crafting inputs that cause the computation to fail while still satisfying the circuit constraints, thereby producing bogus proofs for traces that never actually occurred. Our evaluation shows that such abort vulnerabilities account for nearly 50% of the real-world bugs zkFuzz identified.
Second, when detecting over-constrained vulnerabilities, prior work [24] considers only a special case where $\mathcal{C}$ permits an empty set of traces. A complete definition must consider the full set of execution traces produced by $\mathcal{P}$ and those accepted by $\mathcal{C}$. In particular, it must account for intermediate values. Suppose $\mathcal{P}$ produces a trace $(x,z,y)$ for input $x$, but $\mathcal{C}$ accepts only $(x,z^{\prime},y)$ where $z\neq z^{\prime}$. In this case, $\mathcal{C}$ is over-constrained with respect to $\mathcal{P}$, as an honest prover with the valid trace $(x,z,y)$ would be unable to generate a valid proof. If, however, $\mathcal{C}$ accepts both $(x,z,y)$ and $(x,z^{\prime},y)$ (a common optimization in practice), there is no violation, as the prover can produce a valid proof.
3.2 Trace-Constraint Consistency Test
Let $\mathbb{F}_{q}$ be a finite field of a prime order $q$.
Definition 1 (ZK Program).
A ZK program is defined as a pair of a computation and a set of circuit constraints $(\mathcal{P},\mathcal{C})$:
@@ 3.2 Trace-Constraint Consistency Test
  • On abnormal termination, it returns the special value $(\_,\bot)$.
@@
Definition 2 (Execution Trace).
@@
Definition 4 (Constraint Satisfaction Set).
@@
[…]
@@
Definition 5 (Projection).
Let $\{(x_{1},$$z_{1},$$y_{1}),$$\ldots\}$$\subseteq 2^{\mathbb{F}^{n}_{q}\times\mathbb{F}^{k}_{q}\times\mathbb{F}^{m}_{q}}$ be a set of tuples of inputs, intermediates and outputs. Then, we define the projection operator $\Pi_{xy}$ as $\Pi_{xy}(\{(x_{1},z_{1},y_{1}),\ldots\}):=\{(x_{1},y_{1}),\ldots\}$.
@@
Definition 6 (Under-Constrained Circuit).
@@
$\Pi_{xy}(\mathcal{S}(\mathcal{C}))\setminus\Pi_{xy}(\mathcal{T}(\mathcal{P}))\neq\emptyset$ (1)
Intuitively, $\mathcal{C}$ is under-constrained for $\mathcal{P}$ if $\mathcal{C}$ accepts (x,_,y)(x,\_,y) but $\mathcal{P}(x)=(\_,y^{\prime})$ and $y\neq y^{\prime}$ or $\mathcal{P}(x)$ aborts. Projection out the intermediate values here is necessary for the following reason. In contrast to over-constrained vulnerabilities, if $\mathcal{P}(x)=(z,y)$ and $\mathcal{C}$ accepts $(x,z^{\prime},y)$ such that $z\neq z^{\prime}$, it is not an under-constrained vulnerability, since malicious provers still cannot create bogus proofs asserting that $\mathcal{P}(x)$ yields an output $y^{\prime}\neq y$. Similarly, even if $\mathcal{C}$ accepts both $(x,z,y)$ and $(x,z^{\prime},y)$, adversaries cannot forge proofs with outputs differing from $y$.
This definition naturally takes care of abnormal termination of the computation. Any input $x$ that causes $\mathcal{P}$ to abort will not appear in the trace set $\mathcal{T}(\mathcal{P})$. If $\mathcal{S}(\mathcal{C})$ includes any tuple of the form (x,_,_)(x,\_,\_), then the circuit is under-constrained, as it accepts an input that the computation cannot successfully process.
Definition 7 (Over-Constrained Circuit).
@@
$\mathcal{T}(\mathcal{P})\setminus\mathcal{S}(\mathcal{C})\neq\emptyset$ (2)
@@
We now combine Eq. 1 and Eq. 2 to test the consistency between the trace and the constraint satisfaction sets:
Definition 8 (Trace-Constraint Consistency Test).
@@
$\Pi_{xy}(\mathcal{S}(\mathcal{C}))\setminus\Pi_{xy}(\mathcal{T}(\mathcal{P}))=\emptyset\hskip 5.0pt\land\hskip 5.0pt\mathcal{T}(\mathcal{P})\setminus\mathcal{S}(\mathcal{C})=\emptyset$ (3)
If Eq. 3 holds, we say that $\langle\mathcal{P},\mathcal{C}\rangle$ is well-constrained, which is equivalent to $\mathcal{T}(\mathcal{P})\subseteq\mathcal{S}(\mathcal{C})$ $\land$ $\Pi_{xy}(\mathcal{T}(\mathcal{P}))=\Pi_{xy}(\mathcal{S}(\mathcal{C}))$, meaning that all execution traces satisfy the constraints, and the set of input-output pairs is identical between the execution traces and the constraint satisfaction set.
@@
a) Deciding if $\mathcal{C}$ is under-constrained for $\mathcal{P}$ is NP-complete.
b) Deciding if $\mathcal{C}$ is over-constrained for $\mathcal{P}$ is NP-complete.
c) TCCT is co-NP-complete.
Proof in Appendix A is based on the reduction from the Boolean Satisfiability problem and the tautology problem.
Comparison to Prior Definitions. Tab. I compares TCCT with prior definitions. For under-constrained bugs, prior work detect only non-deterministic behaviors [18, 25], and fail to capture those caused by computation aborts. In addition, they yield false positives when the expected behavior of the witness generation program is non-deterministic, while our TCCT can properly handle those cases (see Appendix C for a detailed discussion). For over-constrained bugs, prior work considers only a special case where the constraint satisfaction set is empty [24]. In contrast, TCCT is more general and complete, capturing all bugs detectable by existing models, as well as additional bugs that prior definitions overlook. A key advantage of TCCT is its DSL-independence: even though different DSLs may offer varied primitives or syntactical sugar for specifying constraints and computations in ZK circuits, TCCT ’s definitions consistently apply across all of them.
TABLE I: Comparison of formal definitions of ZK vulnerabilities. Our TCCT is more general than existing models.
@@
[24] Same as [18, 25] $\mathcal{S}(\mathcal{C})=\emptyset$
@@
3.3 Vulnerable and Benign Examples
@@
1template Transfer(){
2 signal input fb, tb;
3 signal input amt;
4 signal output fn, tn
5 assert(fb - amt >= 0);
6 fn <== fb - amt;
7 tn <== tb + amt;
@@
Code 4: Under-constrained bug from [43] undetectable by any existing formulations, only captured by our TCCT.
@@
7 x*y === 0;
8}
Code 5: Benign circuit from [44] incorrectly flagged by prior static tools. Its trace and constraint satisfaction sets are shown in Tab. II.
Code 4 shows a real-world under-constrained circuit (slightly modified for clarity) caught by zkFuzz from zvm [43]. it had previously gone undetected until our discovery, later confirmed by the developers. The circuit takes as inputs the sender’s balance fb, the recipient’s balance tb, and the transfer amount amt, and produces updated balances fn and tn. This code tries to ensure that fb is greater than amt via assert. (line 5). Circom’s assert, however, performs runtime checks only without adding any constraint. As a result, the computation aborts when fb is higher than amt, the constraints remain satisfiable for any combination of fb and amt. No prior formal tools can detect this bug because this circuit is always deterministic. Existing static tools [13, 20] also do not cover this misuse of assert. To correct this circuit, one must introduce a genuine constraint enforcing fb > amt—for example, by using the LessThan component from [44]. §6.3 also presents a similar bug involving the misuse of assert, which is not detected by any existing tools.
Code 5 shows a benign circuit template from Circom’s de facto standard library, circomlib [44], which returns one if the input $x$ is zero and zero otherwise. Line 6 introduces an intermediate value $z$ using a weak assignment, which does not add any constraint since it is not a quadratic expression. Line 7 assigns the output $y$ and introduces a constraint relating $x$, $z$, and $y$. Line 8 adds both a runtime assertion
TABLE II: Trace and Constraint Satisfaction Sets of IsZero in Code 5 ($q=3$)
@@
and an equality constraint. The full execution trace and constraint satisfaction sets are shown in Tab. II. Notably, $\mathcal{T}(\mathcal{P})$ and $\mathcal{S}(\mathcal{C})$ are not identical: $\mathcal{S}(\mathcal{C})$ includes two additional tuples. However, TCCT correctly handles this case. After projecting out intermediates, the sets of $(x,y)$ pairs are identical, indicating no under-constrained bug. While $\mathcal{S}(\mathcal{C})$ includes three different intermediate values, $z=0,1,2$, for the case where $x=0$ and $y=1$, it includes the valid tuple $(0,0,1)$ from $\mathcal{T}(\mathcal{P})$, so there is no over-constrained bug either. Despite being benign, this circuit is incorrectly flagged by prior static analyzers, which rely on pattern matching rather than a precise semantic definition of ZK bugs.
We refer the readers to Appendix B for more examples, including under-constrained circuits caused by insecure use of the LessThan template, a bug that only zkFuzz can detect, and over-constrained circuits.
4 zkFuzz: Design
@@
4.1 Joint Program Mutation & Input Generation
zkFuzz uses both program mutation and input fuzzing for detecting bugs. Although program mutation is widely used in mutation testing to assess test case effectiveness [45], our objective differs: we leverage mutation specifically to generate alternative traces that continue to satisfy the constraint, exposing undesired behaviors in ZK circuits.
Algo. 1 describes the key steps that examine both $\mathcal{P}$ and its associated constraints $\mathcal{C}$ by using two testing modes:
@@ 4.1 Joint Program Mutation & Input Generation
Algorithm 1 Core Logic of zkFuzz
@@
For simplicity, Algo. 1 shows one input and one program mutant sampled per iteration. Both input sampling and program mutation can be biased toward targets that are more likely to expose vulnerabilities. The algorithm also requires minimal bookkeeping across iterations: once program mutants are generated, the current iteration operates independently of the previous ones.
Figure 2: Basic workflow of zkFuzz. This fuzzing framework systematically mutates program logic and feeds artificially generated input data to the original and mutated programs to catch inconsistencies between the program and the constraint. The error-based fitness function is utilized to steer the selection and crossover of mutants. Several target selectors are also applied for input generation and program mutation to guide the search.
4.2 Guided Search